what is the dimensional formula of  universal garvitational constant?
 

Grade:12th pass

1 Answers

Jobin C J
6 years ago
F=GMM/R^2There fore G=FR^2 / GMMi.e. D.F OF G (Universal Gravitational Constant) = M^-1 L^3 T^-2(M raise to -1, L raise to 3, T raise to -2)